OpenStudy (anonymous):

A circle is centered at (2, -5) and has radius 3. Write its equation in standard form. A) (x + 2)2 + (y - 5)2 = 3 B) (x - 2)2 + (y + 5)2 = 3 C) (x + 2)2 + (y - 5)2 = 9 D) (x - 2)2 + (y + 5)2 = 9

OpenStudy (mathstudent55):

Equation of circle with center \((h, k)\) and radius \(r\). \((x - h)^2 + (y - k)^2 = r^2\)

OpenStudy (mathstudent55):

Replace x with 2, k with -5, and r with 3. What do you get?

OpenStudy (anonymous):

(x-2)^2+(y+5)^2=9

OpenStudy (kayders1997):

That's right :)